OpenStudy (anonymous):

g(x)=-(x+4)^2+7 . i need x intercepts

OpenStudy (anonymous):

set g(x) = to 0 and solve for x. \[-(x+4)^2 + 7 = 0\] You could also solve graphically by plugging in the equation in your calculator and finding where the graph crosses the x-axis. You could also solve using the quadratic formula or by completing the square.

OpenStudy (anonymous):

\[(x+4)^2=7\\ x+4=\pm\sqrt7\\ x=-4\pm\sqrt7\]